OpenStudy (anonymous):

Solve. x – 6 = –14 (Points : 1) –20 20 8 –8

OpenStudy (texaschic101):

step by step....remember that whatever you do on one side of the equal sign, you have to do on the other side. x - 6 = -14 --- add 6 to both sides x - 6 + 6 = -14 + 6 -- combine like terms x = - 8 any questions ?

OpenStudy (anonymous):

thank you so so much any more?

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

Solve. 3 = –(–y + 6) (Points : 1) –9 –3 3 9

OpenStudy (anonymous):

and then three more than I'm done, if you feel like helping me I would be so so grateful but if not I understand

OpenStudy (texaschic101):

3 = - (-y + 6) --- distribute through the parenthesis 3 = y - 6 --- add 6 to both sides 3 + 6 = y - 6 + 6 -- combine like terms 9 = y

OpenStudy (anonymous):

just in case you seen the question mark diamonds

OpenStudy (texaschic101):

I saw them...when I see them, I just refresh the page and their gone...I can then see all the numbers

OpenStudy (texaschic101):

whats the next one

OpenStudy (texaschic101):

you see my answer up above ^^

OpenStudy (anonymous):

i dont see the answer

OpenStudy (texaschic101):

I posted it up there....answer is y = 9

OpenStudy (texaschic101):

got more ?

OpenStudy (anonymous):

thank you sorry had to refresh my page I see it and yes

OpenStudy (texaschic101):

ask away

OpenStudy (anonymous):

Solve for x. |x| – 5 = 7 (Points : 1) x = –12 or 12 x = 12 x = –2 or 2 x = 2

OpenStudy (anonymous):

A?

OpenStudy (texaschic101):

| x | - 5 = 7 --- add 5 to both sides | x | = 7 + 5 | x | = 12 x = 12 and x = -12...yep, its A

OpenStudy (anonymous):

Serena parachutes from a plane at a height of 20,000 ft above sea level. Several minutes later she notes from an instrument on her wrist that she had dropped 15,000 ft. How far above sea level is she now? (Points : 1) 5000 –5000 10,000 –15,000

OpenStudy (anonymous):

after this its one more then I'm DONE :D

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

I think I am just trying to make this one harder then what it is... I think it is just 5000 ft.

OpenStudy (anonymous):

its okay if you dont know this one just please don't leave

OpenStudy (texaschic101):

ok....what else ya got

OpenStudy (anonymous):

A i think

OpenStudy (texaschic101):

I think so too

OpenStudy (anonymous):

laast one ...

OpenStudy (texaschic101):

I am ready...lol

OpenStudy (anonymous):

Given that the variable t represents the temperature at 6 p.m., which equation would best model the word problem below? The temperature at the summit of Mt Washington dropped 19°F between 6 p.m. and 10 p.m. If the temperature at 10 p.m. was –12°F, what was the temperature at 6 p.m.? (Points : 1) 12 – 19 = t 19 + 12 = t t – 19 = –12 19 – t = –12

OpenStudy (texaschic101):

lots of diamonds..let me refresh

OpenStudy (texaschic101):

t - 19 = -12

OpenStudy (anonymous):

okay or do you want me to screenshot?

OpenStudy (texaschic101):

already done.

OpenStudy (anonymous):

thank you are you sure on all of these (just making sure NOT at all saying your wrong just want to make sure)

OpenStudy (anonymous):

yes, I'm done and thank you thank you so so much

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

I, myself, would put these answers on a test that counted for 100% of my grade

OpenStudy (texaschic101):

I am glad I could help :)

OpenStudy (anonymous):

:)